[clas_members] [New Logentry] Daily RC summary

pierrec at jlab.org pierrec at jlab.org
Fri Mar 10 08:50:02 EST 2023

Logentry Text:
RC summary 10th of March

Day (9th of March):
Quiet first half of the shift, beam went on and off as usual. At 14h it was decided to change the target at 15h. Target swap started at 15h as planned after Radcon inspection. Rafo and Sergey took the opportunity to install the DAQ for the muRWell at the same time.
ABU 4.5h

Swing (9th of March):
Target was swapped to carbon, and data collection restarted at 17h after a 2h interruption. 
Despite a usual rich recovery procedure and couple of Roc reboot, the shift was qualified as “boring” by the expert on shift. I personally like this kind of boring shift
ABU 5.1h

Owl (10th of March)
Still taking data on C.
Expert called MCC to verify the beam position at 2H01 as it was fluctuating outside of limits. MCC did the job to stabilize this.
At 7h, MVT sector 2 L1 tripped three times in 30 minutes. The expert power cycled it. It appears to fix it for now. We will call the expert if this occurs again
Overall, another quiet shift of data taking
ABU 5.8h (Nice!)

Plan for the day: Take data on C until 15h or until any long downtime. Work planned on RF this morning might lead to a long downtime. The crew chief will let us know if/when it happens. Radcon is informed that we will need them today. We will do the target change accordingly. 
Longer run plan: Take data on NH3- until Sunday afternoon when we will switch to empty target until Monday configuration change.

PS: There are homemade muffins on the snack table in the CH


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4146710
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20230310/ab88b12d/attachment.html>

More information about the clas_members mailing list